I do not know how to find the x-intercept and the vertex of a quadratic equation. For example I have this problem: y=X^2+8x+8 and I do not know how to find the x-intercept and the vertex. This is a worksheet that my teacher hand wrote.
Answer by scott8148(6628)

when the parabola crosses the x-axis (the x-intercept), the value of y is zero
so to find the x-intercepts, substitute zero for y
0=x^2+8x+8 ___ is not factorable ___ use the quadratic formula to find the roots (x-intercepts)

the vertex is on the axis of symmetry
the equation for the axis of symmetry is x=-b/(2a)
in this case, x=-8/2 ___ or x=-4
substitute to find the y value
y=(-4)^2+8(-4)+8 ___ y=16-32+8 ___ y=-8
so the vertex is (-4,-8)